An open cubed shaped box had a volume of 8x3 what is the total surface area
Answer by stanbon(75887)

An open cubed shaped box had a volume of 8x3 what is the total surface area
------
Since Volume of a cube = side^3
-----
side^3 = (8x^3) = (2x)^3
Therefore side = 2x.
------
Since Surface Area = 6*side^2
----
Your Problem:
SA = 6(2x)^2 = 24x^2
========================
Cheers,
Stan H.
========================